OpenStudy (anonymous):

Solve the radical equation √18x-9=x+4

myininaya (myininaya):

\[\sqrt{18}x-x=4+9\] \[x(1-\sqrt{18})=13\] \[x=\frac{13}{1-\sqrt{18}}\] \[x=\frac{13(1+\sqrt{18})}{1-18}=\frac{13+13\sqrt{18}}{-17}\]

myininaya (myininaya):

but somehow i think you meant to write \[\sqrt{18x-9}=x+4\]

OpenStudy (anonymous):

Yes I did.

myininaya (myininaya):

square both sides

myininaya (myininaya):

\[(\sqrt{18x-9})^2=(x+4)^2\] \[18x-9=(x+4)(x+4)\] \[18x-9=x^2+4x+4x+16\] \[18x-9=x^2+8x+16\] \[18x-18x-9+9=x^2+8x-18x+16+9\] \[0=x^2-10x+25\]

myininaya (myininaya):

does this help?

myininaya (myininaya):

0=(x-5)(x-5) x=5

OpenStudy (anonymous):

Yes.